Suppose you have three resistors of 20 Ω, 50 Ω and 100 Ω. What minimum and maximum resistance can you obtain from these resistors?

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. Max. R = (20 + 50 + 100) Ω = 170 Ω Min R = 1/(1/20 + 1/50 + 1/100) = 100/8 = 12.5 Ω.